(3-z)(4z^3+6x+2)= (4z^3+6x+2) ganna be (64z+6x+2) ok now 6x+2= 8x+64z= 72xz.
now 3-z = 3-1 coz x=-1 and it's = 2 X 72xz= 144xz

Found 2 solutions by checkley77, nknkhotpinkgirl:
Answer by checkley77(12844)

(3-z)(4z^3+6x+2)
12z^3+18x+6-3z^4-6xz-2z
-3z^4+12z^3-2z-6xz+18x+6

Answer by nknkhotpinkgirl(1)


(3-z)(4z^3+6x+2)= (4z^3+6x+2) ganna be (64z+6x+2) ok now 6x+2= 8x+64z= 72xz.
now 3-z = 3-1 coz x=-1 and it's = 2 X 72xz= 144xz..
my answer is 144xz